If the air temperature is the same as the temperature of your skin (about 30.0 oC), your body cannot get rid of heat by transferring it to the air. In that case, it gets rid of the heat by evaporating water (sweat). During bicycling, a typical 70.0kg person's body produces energy at a rate of about 53.6 W due to metabolism, 80.0% of which is converted to heat. How many kilograms of water must the person's body evaporate in an hour to get rid of this heat? The heat of vaporization of water at body temperature is 4.0 ×106 J/kg.
Answer #1

Heat produced by body

Q=MLv

Since

Q=Pt=(53.6*0.8)(1*3600)=154368 J

Therefore

M=Q/Lv =154368/(4*106)

M=0.0386 kg